how do i do "e" the natural base???
e^5 e with the little 5 on top like a square root?

Answer by stanbon(75887)

You probably should have a calculator to do this>
e^5 = e*e*e*e*e= 148.41.....
The value of e is 2.718281828459....
Cheers,
Stan H.